We're back.
Consecutive Intergers - Write the equation and solve.
The problem: Twice the greater of two consecutive odd integers is 13 less than three times the lesser. What are the integers.

Answer by scott8148(6628)

Welcome home...

2(x+2)+13=3x ... 2x+4+13=3x ... 17=x ... 19=x+2